Find h'(3) if h (x)=7x + 4/(x^2-1)^(1/3) .
Answer by Boreal(15235)

I'm assuming h(x)=7x + 4/(x^2-1)^(1/3), two terms.
derivative= 7+4/(x^2-1)^(1/3)=4(x^2-1)^(-1/3)
=7 (-1/3)*4(x^2-1)^(*-4/3)*2x
when x=3,
=7-(8/3)x*(8)^(-4/3), but 8^(-4/3) is 1/8*(4/3) or 1/16; so -(8/3)*3*(1/16)=-1/2
=6.5